# Improvement of Spiking Neural Network with Bit Planes and Color Models

### v2

@@ VII Conclusion
In conclusion, our research introduces a novel coding method for SNN that leverages bit planes derived from various color models of input image data. This approach prove to enhances the computational accuracy of SNN without increasing the model size. Through extensive experimental validation, we have demonstrated the effectiveness of our coding strategy across multiple computer vision tasks, achieving notable performance gains compared to conventional methods. This study is pioneering in its application of color models within the context of spiking neural networks, and we believe that our findings open new avenues for the development of more efficient and effective SNN models. Our work sets a foundation for future research and applications, potentially advancing the field of computational neuroscience and artificial intelligence.
Appendix A CMY or CMYK?
In the initial stages of our experimental setup and algorithm design, we encountered challenges in selecting the appropriate device-dependent color model for our experiments. Specifically, we debated whether to use the CMY or CMYK color model. Each has advantages in different applications, but their compatibility with the neural networks used in our research required careful consideration. In image processing and printing, the CMYK model was developed to improve print quality by reducing ink usage in darker regions. This is achieved by adding a K component, which reduces the CMY values proportionally, leading to enhanced contrast and more economical ink usage in printed images.
However, when adapting this approach for digital image processing, particularly for direct usage on digital devices, the inclusion of a K channel becomes less beneficial. Unlike printed images, digital displays do not require ink density adjustments, making the K channel unnecessary for accuracy. For our experiment, this would introduced dimensional conflicts in tested SNN architectures and coding schemes since they are previously optimized for three-channel data inputs (RGB), including a fourth channel would require additional modifications to model architecture, potentially affecting performance and comparability.
Furthermore, in our image processing pipeline, we observed that converting images to the CMYK color model using standard image processing libraries (e.g., PIL[55]) did not apply undercolor removal during conversion to CMYK model, resulting in a K channel of zero values. This posed a unique problem in our training setup: a zeroed K channel in the input data could nullify both the input signal and its gradients when surrogate gradients are used for backpropagation in SNNs. Consequently, this would undermine the learning process by reducing the effectiveness of gradient-based updates, complicating model training and ultimately impacting the accuracy of our results.
Appendix B Tranning progession details
This section contain the training progression graph for each variant of experiments in Section V-B3 and Section V-B1.
It is noteworthy from Figure 4 and 5 that performance fluctuations were observed in both traditional coding and our coding method. We hypothesize that these fluctuations might related to inherent error rate during sampling process in Poisson-distributed spike trains[3].
Another hypothesis is that although bit planes may possess a certain degree of spatial information, they can also encompass noise on low order bit planes that interferes with the training process or have little to none distinguishable features (which might explain how on low noises single-channel image dataset such as MNIST, the performance of only using bit planes coding remain competitive to traditional rate coding with less amount of time step). Consequently, the selective identification of bit planes that retain valuable spatial information continues to pose a significant challenge.
(a) MNIST
(b) Fashion-MNIST
(c) KMNIST
Figure 3: Comparison in accuracy progression and loss progression between coding methods on different low-resolution gray-scale image classification datasets.
(a) CIFAR10[26]
(b) CIFAR100[26]
(c) Caltech101[48]
(d) Caltech256[49]
(e) EuroSAT[50]
(f) Imagenette[51]
(g) Food101[52]
(h) GTSRB[53]
Figure 4: Comparison in loss progression of traditional rate coding with our method on image classification datasets. On every sub-figure, loss details of every color models is shown on the left and average result of all models on the right with maximum loss difference denoted on top of figure.
(a) CIFAR10[26]
(b) CIFAR100[26]
(c) Caltech101[48]
(d) Caltech256[49]
(e) EuroSAT[50]
(f) Imagenette[51]
(g) Food101[52]
(h) GTSRB[53]
Figure 5: Comparison in accuracy progression of traditional rate coding with our method on image classification datasets. On every sub-figure, accuracy details of every color models is shown on the left and average result of all models on the right with maximum accuracy difference denoted on top of figure.

### v3

@@ VII Conclusion
In this paper, we have introduced a novel coding method for SNN that leverages bit planes of input image data. This approach proved to enhance the computational accuracy of SNN without increasing the model size. Through providing theoretically justified proof and extensive experimental validation, we have demonstrated the effectiveness of our coding strategy with multiple well known coding scheme across different grayscale and color datasets. This study is pioneering in the investigation of bit planes and color models within the context of SNN. We expect that the findings in this paper open new avenues for the development of more efficient and effective SNN models. In future work, we will further investigate the impacts of bit plane coding and color models in different application domains, such as medical analysis and object detection for traffic smart cameras. In addition, the proposed method will be implemented and tested on practical hardware platforms.